OpenStudy (anonymous):

28. Bill rides a bus for 2.5 hours to visit a friend. If the bus travels at about 60 to 63 miles per hour, How far away does bill's friend live?

OpenStudy (anonymous):

D=rt or Distance = (rate)(time) \[D=\frac{ 60miles }{ 1hour }\frac{ 5hours }{ 2 } \\or\\ D=\frac{ 60miles }{ 1hour }\frac{ 2.5hours }{ 1 }\] The unit hours cancels out and leaves us with the unit of distance (miles). So, (60*2.5)miles=150miles or (60*5)miles=300miles (300/2)miles=150miles